# Instruction-Based Molecular Graph Generation with Unified Text-Graph Diffusion Model

### v1

@@ A. Analysis of FCD
Since this metric compares the distance between molecule distributions, it can be misled by molecules generated from other instructions. For example, FCD is equivariant under reordering, meaning it cannot distinguish errors where generated molecules and instructions are exactly cross-matched. For instance, the instruction $S_{1}$ match the molecule ‘CCO’, and instruction $S_{2}$ match the molecule ‘[He]’. If we generate ‘CC’ with instruction $S_{1}$ and ‘[He]’ with instruction $S_{2}$, there’s no doubt that FCD([‘CCO’, ‘[He]’], [‘CC’, ‘[He]’]) ¿ 0. However, If we generate ‘[He]’ with instruction $S_{1}$ and ‘CCO’ with instruction $S_{2}$, which is a totally irrelavant results, we’ll got an unexpected result FCD([‘CCO’, ‘[He]’], [‘[He]’, ‘CCO’]) = 0.
@@
B. Proof of required property
The edge attribution must maintain symmetry. This is conducted since the start of embedding is symmetrical as they have same categories. And therefore the edge attribution maintains symmetry in each layer.
Another required property for graph generation model is to ensure that the node’s attribution in the graph is equivariant to reorderings, meaning that the matrices under reordering can represent the same graph. This property requires no positional embedding in the denoising network. However, we empirically found the importance of positional embedding for better performance.
@@ B. Proof of required property
Therefore, the model is a combination of permutation equivariant blocks.
Proof 3
@@
Following the proof in DiGress, with the lemma(Xu et al. 2022): if a distribution $p(G_{T})$ is invariant to the action of a group G and the transition probabilities $p(G_{t-1}|G_{t})$ are equivariant, them $p(G_{0})$is invariant to the action of G. We apply this result to the special case of permutations:
@@
C. Details for evaluation metrics
Since the model generates a matrix of nodes and edges, although ultimately a SMILES will be parsed based on the matrix, it is not directly generated, which can easily cause the problem of different strings for same graph. Therefore, although molecules can be represented by biological sequence structures, and previous models have also established evaluation indicators from the level of string similarity, this article does not involve these indicators in comparison, including NLP indicators such as BLEU(Papineni et al. 2002), Levenshtein(Miller, Vandome, and McBrewster 2009).
@@ C. Details for evaluation metrics
  • FTS: We employ three fingerprint metrics: MACCS FTS, RDK FTS, and Morgan FTS, where FTS stands for fingerprint Tanimoto similarity. MACCS (Durant et al. 2002), RDK(Schneider, Sayle, and Landrum 2015) and Morgan(Rogers and Hahn 2010). The fingerprints of two molecules are compared using Tanimoto similarity (also known as Jaccard index), and the average similarity over the evaluation dataset is reported. We use RDKIT toolkit.
  • FCD score (Fréchet chemnet distance)(Preuer et al. 2018): Measure molecular similarity based on a pre-trained ”ChemNet” bioinformatics network. We use fcd 1.1 in python.
D. Details for experiments
@@
The specific data is listed in Table 11, 12, 13
@@
Here we list the prompt for the three datasets: For CHEBI-20 dataset, We give its instruction form as : ”’[molecule description]’ is the description of molecule:” For the two editing dataset, since there’s task instruction in datasets, we don’t use any additional instructions. Here’s an example in Retrosynthesis dataset: ”Please suggest potential reactants for the given product.”; Here’s an example in forward reaction prediction dataset: ”With the provided reactants and reagents, propose a potential product.”
@@
E. Source of datasets

### v2

@@ Appendix A Analysis of FCD
Since this metric compares the distance between molecule distributions, it can be misled by molecules generated from other instructions. For example, FCD is equivariant under reordering, meaning it cannot distinguish errors where generated molecules and instructions are exactly cross-matched. For instance, the instruction $S_{1}$ match the molecule ‘CCO’, and instruction $S_{2}$ match the molecule ‘[He]’. If we generate ‘CC’ with instruction $S_{1}$ and ‘[He]’ with instruction $S_{2}$, there’s no doubt that FCD ([‘CCO’, ‘[He]’], [‘CC’, ‘[He]’]) > 0. However, If we generate ‘[He]’ with instruction $S_{1}$ and ‘CCO’ with instruction $S_{2}$, which is a totally irrelavant results, we’ll got an unexpected result FCD ([‘CCO’, ‘[He]’], [‘[He]’, ‘CCO’]) = 0.
@@
Appendix B Proof of equivarient property
The graph generation model needs to ensure that the node’s attribution in the graph is equivariant to reorderings, meaning that the matrices under reordering can represent the same graph. This property requires no positional embedding in the denoising network. However, we empirically found the importance of positional embedding for better performance.
@@ Appendix B Proof of equivarient property
Therefore, the model is a combination of permutation equivariant blocks. ∎
Proof.
@@
Following the proof in DiGress, with the lemma: if a distribution $p(G_{T})$ is invariant to the action of a group G and the transition probabilities $p(G_{t-1}|G_{t})$ are equivariant, them $p(G_{0})$is invariant to the action of G. We apply this result to the special case of permutations:
@@
Appendix C Details for evaluation metrics
Since the model generates a matrix of nodes and edges, although ultimately a SMILES will be parsed based on the matrix, it is not directly generated, which can easily cause the problem of different strings for same graph. Therefore, although molecules can be represented by biological sequence structures, and previous models have also established evaluation indicators from the level of string similarity, this article does not involve these indicators in comparison, including NLP indicators such as BLEU, Levenshtein.
@@ Appendix C Details for evaluation metrics
  • FTS: We employ three fingerprint metrics: MACCS FTS, RDK FTS, and Morgan FTS, where FTS stands for fingerprint Tanimoto similarity. MACCS [10], RDK[44] and Morgan[42]. The fingerprints of two molecules are compared using Tanimoto similarity (also known as Jaccard index), and the average similarity over the evaluation dataset is reported. We use RDKIT toolkit.
  • FCD score (Fréchet chemnet distance): Measure molecular similarity based on a pre-trained "ChemNet" bioinformatics network. We use fcd 1.1 in python.
For better evaluation, we also list some metrics which is not discussed in previous baseline. Here’re the information and results:
  • Maximum Common Edge Substructure (MCES) : This metric finds the largest subgraph (in terms of the number of edges) that is common to two or more given graphs. We used the rdRascalMCES function in RDKit (Python) to measure MCES, with the default threshold set to 0.7. If either of the Johnson similarity estimates falls below this threshold, we consider as invalid for MCES.
The results of MCES in CHEBI-20 is shown in Table 12.
Method MCES
BioT5 (reproduce) 0.8891
gpt4-2024-04-09 (10 shot+BM25) 0.9060
TGM-DLM 0.9036
UTGDiff (our) 0.9166
Table 12: MCES results in different models
Appendix D Details for experiments
@@
The specific data is listed in Table 14, 15, 16
@@
Here we list the prompt for the three datasets: For CHEBI-20 dataset, We give its instruction form as : "’[molecule description]’ is the description of molecule:" For the two editing dataset, since there’s task instruction in datasets, we don’t use any additional instructions. Here’s an example in Retrosynthesis dataset: "Please suggest potential reactants for the given product."; Here’s an example in forward reaction prediction dataset: "With the provided reactants and reagents, propose a potential product."
Also, here we list the prompt for large language model baselines. To limit the Generalist LLM to generate the desired SMILES, we use below prompts for both chatgpt and gpt-4 baseline:
[…]
@@
Appendix E Algorithm Pseudo Code
The pseudo code of the training and sampling algorithm is shown in 1 and 2.
Algorithm 1 Training Algorithm
0: A graph $G=(V,E)$ with description $S$
1:repeat
2:  Sample $t\sim U(1,\cdot,T)$
3:  Sample $G^{t}\sim V\bar{Q}_{V}^{t}\times E\bar{Q}_{E}^{t}$
4:  $\hat{p}(S),\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
5:  loss = $l_{CE}(\hat{p}(E),E)+l_{CE}(\hat{p}(V),V)+l_{CE}(\hat{p}(S),S)$
6:  optimizer.step(loss)
7:until converged
Algorithm 2 Sampling Algorithm
[…]
2:for $t=T$ to 11 do
3:  $\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
7:end for
8:return  $G^{0}$
Appendix F Potential Risks
Potential risks associated with molecular generation models include the inadvertent creation of toxic or unstable compounds, which could pose significant safety hazards. Additionally, biases in the training data could be learned by the model, leading to unintended consequences. Another risk is the possibility of the model being used to discover harmful molecules instead of beneficial ones. These concerns highlight the necessity for stringent regulatory frameworks and ethical guidelines to govern the use of these models.
Appendix G Source of codes and datasets
Our Code is avaliable at: https://github.com/ran1812/UTGDiff